In an electric shaver, the blade moves back and forth over a distance of 2.0mm in simple harmonic
motion, with a frequency 120Hz. Find (a) the amplitude, (b) the maximum blade speed and (c) the
magnitude of the maximum acceleration of blade.
1
Expert's answer
2020-07-21T13:49:09-0400

(a) "A=L\/2=2\/2=1mm"


(b) "v=\omega A=2\pi fA=2\cdot3.14\cdot120\cdot0.001=0.75m/s"


(c) "a=\omega^2A=(2\pi f)^2A=(2\cdot3.14\cdot 120)^2\cdot 0.001=568m/s^2"